1.)The grams of water that need to be added to 17.0g of BaCl2 to prepare a 14.0% Bacl2 solution.

2.) You have a solution that is 60.0% HCl (MW 36.5g/mol). What volume, in mL, of this solution contains 0.500 mol of HCl? The density of the solution 1.28g/ML.
Answer #1

ILE = 11 X 100. wel Seluten L 景 .hot x to xd м= 60 x 10 x 1.28 36:5 s(tit) 0.500 M21-04 Nolurne inmı= o-0838x1000 = 938 m)